### v1

# Groups with et0l co-word problem

Abstract
We study groups whose co-word problems are ET0L languages, which we call coET0L groups, using an automaton based model due to van Leeuwen, and recently studied by Bishop and Elder. In particular we prove a number of closure results for for the class of groups with co-words problems in a subclass of ‘special’ ET0L languages; that class of groups contains all groups that we know at the time of writing to be co-ET0L, including all groups that were proved by Holt and Röver to be stack groups, and hence co-indexed. It includes virtually free groups, bounded automata groups, and the Higman-Thompson groups, together with groups constructed from those using finitely generated subgroups, finite extension, free and direct products, and by taking the restricted standard wreath product of a co-et0l group by a finitely generated virtually free top group.
2010 AMS Subject classification: 20F10, 68Q42, 68Q45.
Keywords: ET0L language, check stack pushdown automaton, indexed language, word problem, co-word problem.
1 Introduction
A finitely generated group is said to be co-et0l if the complement of its word problem (i.e. its co-word problem) lies in the class of et0l languages. Our aim in this article is to prove that certain types of groups are co-et0l, and to establish some closure results for this class of groups. More precisely we prove these results for the class of groups with co-word problem that is a special et0l language; that class of groups contains all groups that we know at the time of writing to be co-et0l, and it coincides with the class of all groups that are proved in [9] to be co-indexed. In particular, it includes virtually free groups, bounded automata groups, and the Higman-Thompson groups, together with groups constructed from those using finitely generated subgroups, finite extension, free and direct products, and by taking the restricted standard wreath product of a co-et0l group by a finitely generated virtually free top group.
Various classes of formal languages of higher complexity than that of context-free languages have been examined as candidates to define natural complexity classes for certain problems in group theory, since Muller and Schupp’s seminal result [13] characterising finitely generated virtually free groups via their context-free word problems. Indexed languages, corresponding to indexed grammars and nested stack automata, entered the arena with Bridson and Gilman’s result [4] finding combings for the fundamental groups of all compact 3-manifolds, but the subclass (within the indexed languages) of et0l languages started to be studied after results of Ciobanu, Diekert and Elder in [6, 7] showed that such languages contained solution sets to equations over first free and then hyperbolic groups. Holt and Röver had proved in [9] that all finitely generated bounded automata groups (and so in particular Grigorchuk’s group of intermediate growth and the Gupta-Sidki group) are co-indexed (i.e. they have co-word problems that are indexed languages), as well as the Higman-Thompson groups (later proved in [12] to be co-context free). At the same time they had proved various closure properties for a subset of the family of co-indexed groups, whose members they called stack groups and which contained all groups that they could prove to be co-indexed. Ciobanu and Elder [7] then proved that the Grigorchuk group is in fact co-et0l. Subsequently Bishop and Elder [2] derived the same result for all finitely generated bounded automata groups, using an automata-based characterisation of et0l languages.
In this article, motivated by the characterisation of et0l languages that was used in [2] (and attributed to van Leeuwen [16]) and by closure results for specific co-et0l groups appearing in the thesis work of Al Kohli [1] (it is proved in [1] that the free products ${\mathbb{Z}}^{m}*{\mathbb{Z}}^{n}$ and ${\mathbb{Z}}^{m}*G$, for $G$ virtually free, are co-et0l), we sought to convert the results of [9] from indexed to a restricted class of et0l languages. It had been suggested in [1] that the results of [9] would carry over to the class of groups which we shall refer to here as special co-et0l groups. Specifically, we have converted the restrictions on nested stack automata that were imposed on those recognising the co-word problems of the stack groups of [9], in order to define the classes of special et0l languages, and of special co-et0l groups.
We prove in Theorem 3.14.15.1 that the class of special co-et0l groups contains all virtually free groups, all finitely generated bounded automata groups and the Higman–Thompson groups, extending to special co-et0l the results of [13, 2, 9]. In Theorem 6.1 we prove closure of the class of special co-et0l groups under passing to finitely generated subgroups, finite direct products and finite extensions, and that the property of being special co-et0l is independent of finite generating set. In Theorems 6.2 and 6.3 we prove closure under taking the restricted wreath product with a finitely generated virtually free group, and under taking free products with finitely many free factors.
We are grateful to Murray Elder for some very helpful conversations about the cspda approach to et0l languages, and in particular for the formulation of Property (3) in our definition of a special cspda below.
2 Definitions and assumptions
Our definition of a check stack pushdown stack automaton (which we shall abbreviate as cspda) is basically taken from [2]. Automata with the same name were also introduced by van Leeuven in [16], with a slightly different definition. However, it is proved in [16], and also in an Appendix to [2] that is attached to the unpublished version  [3], that the class of languages that are recognisable by the cspda defined in the respective article is precisely the class of et0l languages, so these definitions are equivalent, and we will use the one from [2].
We define a cspda to be a non-deterministic finite-state automaton with a one-way input tape, with a read write head that has access to both a check stack (with stack alphabet $\Delta$) and a pushdown stack (with stack alphabet $\Gamma$), where access to these two stacks is tied in a very particular way. The execution of a cspda is separated into two stages.
In the first stage the machine is allowed to push to its check stack but not to its pushdown stack, and further, the machine is not allowed to read from its input tape. Since the words over $\Delta$ that can be pushed to the check stack are controlled by a finite state automaton, the language consisting of all such words is regular. We shall refer to this first stage as the initialisation of the cspda.
In the second stage, the machine can no longer alter its check stack, but is allowed to access its pushdown stack and input tape. We restrict the machine’s access to its stacks so that it can only move along its check stack by pushing or popping items to or from its pushdown stack. In particular, every time the machine pushes a value onto the pushdown stack its read write head will move up the check stack, and every time it pops a value off the pushdown stack it will move down the check stack. So, throughout the second stage of the computation, the read write head can see the top symbol of the pushdown stack and the symbol in the correspondiong position on the check stack.
The machine accepts its input word if it enters an accepting state at some stage after reading all of that input.
We call a state of a cspda a fail state if it is non-accepting with no path from it leading to an accepting state, a reading state if it allows only reading moves, and a non-reading state if it allows only non-reading moves. A configuration of the cspda (that is, a triple $(q,\mathbf{y},\mathbf{z})$, consisting of a state and with the contents of the two stacks) is called a fail configuration when $q$ is a fail state, an accepting configuration when $q$ is an accepting state, and a reading configuration when $q$ is a reading state.
It is clear that, by introducing new states if necessary, we can modify any cspda to define one with the same language for which every state is either reading or non-reading, and which has at most one fail state, which is non-reading state and from which no further transitions are possible We shall assume that each cspda we are working with has both of these properties. In addition, we shall need to assume the following.
  1. (0)
    All paths in the cspda lead eventually to either the fail state or to a reading state; that is, there are no infinite paths of non-reading moves.
We shall be restricting our attention to cspda that satisfy three further properties, (1)–(3). These properties for the cspda relate to the additional properties of the nested stack automata that are introduced in [9, Section 7] for the purpose of establishing various closure properties of stack groups. We shall introduce properties (1) and (2) now, then property (3) at the end of this section.
  1. (1)
    The automaton is deterministic upon input. More precisely, after the completion of the first stage and at the beginning of the second stage, the pushdown stack is empty and the automaton is in a reading state (so the read write head is at the bottom of the two stacks), and the operation of the cspda throughout the second stage is deterministic.
  2. (2)
    During the second stage, the read write head never moves above the top of the check stack, and the cspda fails if it attempts to do so.
Note that the Property (2) can be achieved in any cspda without changing the accepted language, by including a padding symbol ${\sf p}$ in the check stack alphabet $\Delta$, and placing an arbitrary number of symbols ${\sf p}$ at the top of the check stack after completing the first stage in the original cspda. In this case, each transition of the original cspda for which a symbol $\gamma$ is read from the top of the pushdown stack while the read write head is above the top of the check stack corresponds to a transition of the modified automaton for which $\gamma$ is read from the top of the pushdown stack and the symbol ${\sf p}$ is read from the check stack.
We assume throughout this article that the two stack alphabets $\Gamma$ and $\Delta$ of any cspda both contain a padding symbol ${\sf p}$, which can be used as above to ensure property (2), but also for other purposes, as we shall see later.
Given a reading configuration $C$ of the automaton, and a word $w$ over the input alphabet, we can unambiguously define $C^{w}$ to be the machine’s configuration as it enters its first reading state after reading all of $w$, or a failing configuration if it enters the fail state without reaching such a reading state. In the second case we say that $C^{w}$ fails. We can assume that the reading states and configurations entered into at the beginning of the second stage form distinguished subsets, which we call the entry states and entry configurations of the automaton; by contrast a state and configuration of the cspda at the beginning of stage 1 are called an initial state and initial configuration. We shall assume that a cspda has a single initial state.
Our third property is that the cspda $\mathcal{M}$ ignores subwords of input words that are not in its language $\mathcal{L}(\mathcal{M})$, or more precisely:
  1. (3)
    Let $C$ be an entry configuration of $\mathcal{M}$, and $u,v\in A^{*}$ with $v\not\in\mathcal{L}(\mathcal{M})$. Then either $C^{uv}=C^{u}$, or $C^{uv}$ fails.
It follows from Property (3) that, if $C$ is an entry configuration and $v\not\in\mathcal{L}(\mathcal{M})$, then either $C^{v}$ is a fail configuration or $C^{v}=C$; so if $C^{w}\neq C$ then $w\in\mathcal{L}(\mathcal{M})$. Hence we can assume that all reading states other than the entry states are accepting states.
We call a cspda that satisfies Properties (0)–(3) a special cspda for $G$, and a language accepted by such a cspda a special et0l language. (But, as we observed earlier, Property (2) does not in itself involve any restriction on this language class.)
In this article, we are interested in cspda accepting co-word problems of finitely generated groups $G$. So we shall assume from now on that the input alphabet of a cspda under consideration consists of a (finite) inverse-closed generating set $A$ of such a group $G$, and that $A\subseteq\Delta$ and $A\subseteq\Gamma$. We denote the word and co-word problems of $G$ with respect to $A$ by $\mathsf{WP}(G,A)$ and $\mathsf{coWP}(G,A)$.
We call a finitely generated group a special co-et0l-group if, for some finite inverse-closed generating set $A$ of $G$, there is a special cspda with input alphabet $A$ that accepts the language $\mathsf{coWP}(G,A)$. We shall prove in Theorem 6.1 below that the property of being a special co-et0l-group does not depend on the choice of the generating set $A$. We note that the Property (3) that is required for the cspda when $G$ is a special co-et0l-group is precisely what was referred to as ignoring the word problem of $G$ in [9]. We also note that every special co-et0l group is a stack group (in the sense of [9]).
The following lemma is useful for avoiding the failure of $C^{uv}$ in the situation described in Property (3). It will be used in the proof of closure of the class of special co-et0l-groups under free products in Theorem 6.3 below.
Lemma 2.1.
Let $G=\langle A\rangle$ be a group, $\mathcal{M}$ a special cspda with $\mathcal{L}(\mathcal{M})=\mathsf{coWP}(G,A)$, $w\in\mathcal{L}(\mathcal{M})$, and $n\in{\mathbb{Z}}$ with $n>0$. Then there is an entry configuration $C$ of $\mathcal{M}$ for which $C^{w}$ is accepting, and such that $C^{v}$ does not fail for any word $v$ over $A$ of length at most $n$.
Proof.
Let $v_{1},v_{2},\ldots,v_{N}$ be an enumeration of all words over $A$ of length at most $n$. Then the word $w^{\prime}:=v_{1}v_{1}^{-1}v_{2}v_{2}^{-1}\cdots v_{N}v_{N}^{-1}w$ lies in $\mathsf{coWP}(G,A)$, and so there is an entry configuration $C$ such that the state of $C^{w^{\prime}}$ is accepting. It follows that for any prefix $w^{\prime\prime}$ of $w^{\prime}$, $C^{w^{\prime\prime}}$ does not fail.
For any $i$ with $0\leq i<N$, it follows from Property (3) of a special cspda (applied with $u=\epsilon$ and $v=v_{1}v_{1}^{-1}\cdots v_{i}v_{i}^{-1}$) that $C^{v_{1}v_{1}^{-1}\cdots v_{i}v_{i}^{-1}}=C$ and hence that $C^{v_{1}v_{1}^{-1}\cdots v_{i}v_{i}^{-1}v_{i+1}}=C^{v_{i+1}}$. So $C^{v_{i+1}}$ does not fail.
Similarly, we deduce from Property (3) that $C^{v_{1}v_{1}^{-1}\cdots v_{N}v_{N}^{-1}}=C$ and hence $C^{w^{\prime}}=C^{w}$. So $C^{w}$ is accepting, as required. ∎
3 Virtually free groups
Theorem 3.1.
Every finitely generated virtually free group is a special co-et0l-group.
We note that the fact that finitely generated virtually free groups are co-et0l-groups follows immediately from the inclusion of the context-free languages in et0l (proved in [11, Theorems 2.4,2.7]) and the fact that virtually free groups have deterministic context-free word problem, and hence context-free co-word problem. But here we prove additionally that these groups are special co-et0l-groups.
Proof.
Let $G$ be finitely generated and virtually free. We shall prove in Theorem 6.1 below that the property of being a special co-et0l-group does not depend on the choice of the finite generating set of the group $G$. We choose a generating set $A=B\cup T^{\pm 1}$ of $G$ in which $B$ is the inverse closure of a free generating set of a free subgroup $F$ of finite index in $G$ and $T$ is a right transversal of $F$ in $G$ with the element in $F$ removed. Then every element of $G$ can be written uniquely as a word $wt$ with $w$ a freely reduced word in $B^{*}$ and $t\in T\cup\{\epsilon\}$, where $\epsilon$ denotes the empty word. It was shown in [13] that the word problem $\mathsf{WP}(G,A)$ is the language of a deterministic pushdown automaton with states corresponding to $T\cup\{\epsilon\}$, that stores the freely reduced word $w$ on its stack. We can amend that to a pushdown automaton $\mathcal{P}_{G}$ that accepts by empty stack, by storing $t$ (if non-empty) on the stack above $w$.
We convert $\mathcal{P}_{G}$ to a cspda by attaching a check stack onto which we write a string ${\sf p}^{k}$ at the start, for some arbitrary integer $k>0$. After the start, the cspda then operates as $\mathcal{P}_{G}$, additionally reporting failure if at any stage the pushdown stack contains more than $k$ cells (and so the top of the pushdown stack is above the top of the check stack). The new automaton accepts if and only if $\mathcal{P}_{G}$ does not accept.
It is straightforward to verify that this cspda is special. Note that failure can only arise as a result of inadequate initialisation of the check stack, and this can be remedied without otherwise affecting its operation by pushing more p symbols onto the check stack. ∎
4 Bounded automata groups
This section is devoted to the proof of the following result, which extends the main result of [2]; its proof is based on the proof of that result and the proof of the corresponding result in  [9] on which it is based.
Theorem 4.1.
Every finitely generated bounded automata group is a special co-et0l-group.
Bounded automata groups were introduced by Sidki in [14]. The best known examples are the Grigorchuk group and the Gupta–Sidki groups, and more recently studied examples include the basilica groups [5]. Our definition below describes them as groups of permutations of sets of infinite strings, which are generated by finite state transducers (i.e. automata with input and output). Finitely generated bounded automata groups were proved to be co-indexed in [9], and then to be co-et0l in [2].
For $d\geq 2$, we consider the set $\Sigma^{*}$ of strings over the finite set $\Sigma$ of $d$ elements, and write $\epsilon$ for the empty string. There is a natural bijection between $\Sigma^{*}$ and the set of vertices of the $d$-regular rooted tree $\mathcal{T}_{\Sigma}$, in which the $d$ edges down from any given vertex are labelled by the elements of $\Sigma$; similarly there is a natural bijection between $\Sigma^{\omega}$ and the set of ends of $\mathcal{T}_{\Sigma}$. We consider the group of all permutations of $\Sigma^{*}$ that induce automorphisms of $\mathcal{T}_{\Sigma}$, which we shall denote here by $\mathcal{A}_{\Sigma}$ (it is denoted by $\mathsf{Aut}(\mathcal{T}_{d})$ in [2]); such a permutation $\alpha$ must preserve ‘levels’ in the tree, and hence for any $k$ it permutes the set of subsets of strings with common prefixes of length $k$; that is the subsets $\mathbf{x}\Sigma^{*}$ of $\Sigma^{*}$ with $\mathbf{x}\in\Sigma^{k}$. In addition it must have the property that, for any $\mathbf{x}\in\Sigma^{*}$, $\mathbf{y}\in\Sigma^{*}\cup\Sigma^{\omega}$, $(\mathbf{x}\mathbf{y})^{\alpha}$ has $\mathbf{x}^{\alpha}$ as a prefix, and the restriction $\alpha|_{\mathbf{x}}$ of $\alpha$ to $\mathbf{x}$ is defined by $(\mathbf{x}\mathbf{y})^{\alpha}=\mathbf{x}^{\alpha}\mathbf{y}^{\alpha|_{\mathbf{x}}}$.
An automorphism $\alpha\in\mathcal{A}_{\Sigma}$ is called an automaton automorphism if there is an associated finite state transducer $\mathcal{N}_{\alpha}$ such that for any $\mathbf{x}\in\Sigma^{*}$ the image $\mathbf{x}^{\alpha}$ of $\mathbf{x}$ is equal to the output string $\mathbf{x}^{\mathcal{N}_{\alpha}}$ of the transducer when $\mathbf{x}$ is input. An automorphism is called bounded if for some $N\in{\mathbb{N}}$ and every $k\in{\mathbb{N}}$ there are at most $N$ strings $\mathbf{x}\in\Sigma^{k}$ with $\alpha|_{\mathbf{x}}\neq 1$. The set of all bounded automaton automorphisms forms a subgroup of $\mathcal{A}_{\Sigma}$, which we shall call $\mathcal{D}_{\Sigma}$. A bounded automata group is defined to be a subgroup of $\mathcal{D}_{\Sigma}$, for some finite set $\Sigma$.
An automorphism $\phi\in\mathcal{A}_{\Sigma}$ is called finitary if, for some $k_{\phi}\in{\mathbb{N}}$ and all $\mathbf{x}$ with $|\mathbf{x}|\geq k_{\phi}$, the restriction $\phi|_{\mathbf{x}}$ is equal to the identity. We refer the reader to [2] for the precise definition of a directed automaton automorphism; such an automorphism $\delta$ has an associated spine, an element of $\Sigma^{\omega}$, with the property that $\delta|_{\mathbf{y}}$ is finitary for any $\mathbf{y}$ that is not a prefix of $\mathsf{spine}(\delta)$. It is proved in [2, Lemma 1] that the spine of any directed automaton automorphism is eventually periodic, that is, of the form $\mathbf{p}\mathbf{q}^{\omega}$, for fixed strings $\mathbf{p},\mathbf{q}\in\Sigma^{*}$, and that $(\mathbf{p}\mathbf{q}^{\omega})^{\delta}=\mathbf{p}^{\prime}\mathbf{q}^{\prime\omega}$, for fixed strings $\mathbf{p}^{\prime},\mathbf{q}^{\prime}\in\Sigma^{*}$ that have the same lengths as $\mathbf{p}$ and $\mathbf{q}$. Furthermore, if $\mathbf{q}=b_{1}\cdots b_{t}$ with $b_{i}\in\Sigma$, then $\delta|_{\mathbf{p}\mathbf{q}^{k}b_{1}\cdots b_{j}}=\delta|_{\mathbf{p}b_{1}\cdots b_{j}}$ for any $k\geq 0$ and $0\leq j<t$.
Our proof of Theorem 4.1 (below) is a modification of the proof of [2], which itself emulates the proof in [9]; We shall describe a construction that is basically the construction of [2], with some modification to fit the constraints that we have imposed on a cspda to make it special. We shall not repeat proofs of properties of the cspda that are already proved in [2].
We note that in [2] bounded automata groups are described as groups of automorphisms of rooted trees, but we prefer to view them as described above, in terms of their associated actions on sets of finite strings as well as the sets of infinite strings that correspond to the ends of the trees. We also note that the entry states in our cspda are analogous to states of the nested stack automaton constructed in [9] that are called start states in that article.
Proof of Theorem 4.1. Let $G$ be a subgroup of $\mathcal{D}_{\Sigma}$ with finite generating set $X$.
It is proved in [9, Proposition 16] that $\mathcal{D}_{\Sigma}$ is generated by the infinite subset of all of its finitary and directed automaton automorphisms. Hence the subgroup $G$ of $\mathcal{D}_{\Sigma}$ is a finitely generated subgroup of a group with finite generating set $Y$, each of whose elements is either finitary or a directed automaton automorphism. It will follow from Theorem 6.1 below that $G$ is a special co-et0l group provided that the subgroup $\langle Y\rangle$ of $\mathcal{D}_{\Sigma}$ is special co-et0l. Hence, as in [2], we may assume that $G$ itself has such a generating set $Y$.
The proof in [2] that $G$ is co-et0l defines a cspda that provides a 3-stage process to verify non-deterministically that a given element $\alpha\in G$ is non-trivial. In Stage 1, an arbitrary string $\mathbf{x}=x_{1}\cdots x_{m}\in\Sigma^{*}$ is put onto the check stack, with $x_{1}$ as top symbol. The image $\mathbf{x}^{\alpha}$ is computed in Stage 2, which is then compared with $\mathbf{x}$ in Stage 3 and, if $\mathbf{x}\neq\mathbf{x}^{\alpha}$, then $\alpha$ is proved non-trivial. If $\alpha$ is non-trivial, then there must be some such word $\mathbf{x}$ with $\mathbf{x}\neq\mathbf{x}^{\alpha}$, and so the language of the cspda is equal to $\mathsf{coWP}(G,Y)$, as required.
Stage 1 of cspda that we define here is the same as in [2], with $x_{1}$ as the top symbol on the check stack. In the construction of [2], $\mathbf{x}$ is put immediately onto the pushdown stack at the end of Stage 1 but, in our modification, to ensure that our cspda is special, we defer copying the contents of the check stack onto the pushdown stack until after we have entered an entry state of the cspda and read the first letter in the input word. This is because Property (1) of a special cspda requires that the pushdown stack be empty when the first input symbol is read.
During the second stage of the process described in [2], the contents of the pushdown stack are modified so that it contains $\mathbf{x}^{\alpha}$ by the end of that stage; successive images of $\mathbf{x}$ under the generators in the string $\alpha$ are computed, as explained below. During the third stage of that process, the contents of the two stacks are compared with each other, to see if they differ.
Our Stage 2 process is similar to that of [2], but it does a little more than in [2]. For, in order to maintain determinism after Stage 1, we do not have a Stage 3, but do our checking within Stage 2; the checking will be explained later. At Stage 2, we have just two reading states. The first is the unique entry state and is non-accepting, and the second is an accepting state; we shall call those two states $q_{1}$ and $q_{2}$ respectively. All non-reading states are non-accepting.
In our description of Stage 2, we denote the current image $\mathbf{x}^{\beta}$ of $\mathbf{x}$ under the prefix $\beta$ of the input word that has been read so far by $\mathbf{y}=y_{1}y_{2}\cdots y_{m}$. We need to explain how we compute the image of $\mathbf{y}$ under finitary and directed automorphisms in Stages 2a and 2b, respectively.
Stage 2a of the process described in [2] computes the image of $\mathbf{y}$ under a finitary element. Since this involves changing only a bounded number of symbols at the top of the pushdown stack, this is easily done deterministically using finite memory, and we shall not give the details here; they can be found in [2].
Stage 2b of the process described in [2] computes the image of $\mathbf{y}$ under a directed automaton automorphism $\delta$. Let $\mathbf{z}=y_{1}\cdots y_{\ell}$ be the maximal prefix of $\mathbf{y}$ that is also a prefix of $\mathsf{spine}(\delta)$. We shall assume that $\ell<m$; the case $\ell=m$ (where $\mathbf{z}=\mathbf{y}$) is an easy modification of the case $\ell<m$. As we saw above, we have $(\mathbf{p}\mathbf{q}^{\omega})^{\delta}=\mathbf{p}^{\prime}\mathbf{q}^{\prime\omega}$, for fixed strings $\mathbf{p}=a_{1}\cdots a_{s},\mathbf{q}=b_{1}\cdots b_{t}$, $\mathbf{p}^{\prime}=a^{\prime}_{1}\cdots a^{\prime}_{s},\mathbf{q}^{\prime}=b^{\prime}_{1}\cdots b^{\prime}_{t}\in\Sigma^{*}$.
Then, either $\ell\leq s$, in which case $\mathbf{z}=a_{1}\cdots a_{\ell}$ and
$\mathbf{y}^{\delta}=a^{\prime}_{1}\cdots a^{\prime}_{\ell}(y_{\ell+1})^{\delta|_{\mathbf{z}}}(y_{\ell+2}\cdots y_{\ell_{\phi}})^{\phi}y_{\ell_{\phi}+1}\cdots y_{m};$
or $\ell>s$ and, for some $k\geq 0$ and $j$ with $0\leq j<t$, we have $\ell=s+kt+j$, $\mathbf{z}=\mathbf{p}\mathbf{q}^{k}b_{1}\cdots b_{j}$ and
$\mathbf{y}^{\delta}=\mathbf{p}^{\prime}\mathbf{q}^{\prime k}b^{\prime}_{1}\cdots b^{\prime}_{j}(y_{\ell+1})^{\delta|_{\mathbf{z}}}(y_{\ell+2}\cdots y_{\ell_{\phi}})^{\phi}y_{\ell_{\phi}+1}\cdots y_{m},$
where in both cases $\phi$ is the finitary automorphism $\delta|_{\mathbf{z}y_{\ell+1}}$ with associated bound $k_{\phi}$ on the length of the prefix of words that it moves, and $\ell_{\phi}:=\min(\ell+1+k_{\phi},m)$.
We need to pop the prefix of $\mathbf{y}$ of length $\ell_{\phi}$ from the stack. Then we compute its image under $\delta$ and then push that (or rather its reverse, which we denote by a superscript $\mathsf{R}$) onto the stack.
The case where $\ell\leq s$ is straightforward. The computation and storage of $y_{\ell+1}^{\delta_{\mathbf{z}}}$ and $(y_{\ell+2},\cdots,y_{\ell_{\phi}})^{\phi}$ need a bounded amount of memory, and so can be done deterministically using non-reading states and transitions of the cspda. Then we push $((y_{\ell+2}\cdots y_{\ell_{\phi}})^{\phi})^{\mathsf{R}}$ onto the pushdown stack followed by $(y_{\ell+1})^{\delta|_{\mathbf{z}}}a^{\prime}_{\ell}\cdots a^{\prime}_{1}$.
The case where $\ell\geq s$ is more complicated. As explained above, it is proved in [2, Lemma 1] that $\delta|_{\mathbf{z}}=\delta|_{\mathbf{p}\mathbf{q}^{k}b_{1}\cdots b_{j}}$ does not depend on $k$, and so $\delta|_{\mathbf{z}}$ and the finitary automorphism $\phi$ depend only on $\mathbf{p}$, $\mathbf{q}$ and $j$. So the computation and storage of the components of $\mathbf{y}^{\delta}$, namely $\mathbf{p}^{\prime}$, $\mathbf{q}^{\prime}$, $y_{\ell+1}^{\delta_{\mathbf{z}}}$ and $(y_{\ell+2},\cdots,y_{\ell_{\phi}})^{\phi}$, is again straightforward, and needs only a bounded amount of memory. As above, we can then push $((y_{\ell+2}\cdots y_{\ell_{\phi}})^{\phi})^{\mathsf{R}}$ onto the pushdown stack followed by $(y_{\ell+1})^{\delta|_{\mathbf{z}}}b^{\prime}_{j}\cdots b^{\prime}_{1}$ But then we need to push $k$ copies of $\mathbf{q}^{\prime\mathsf{R}}$ followed by $\mathbf{p}^{\prime\mathsf{R}}$ without knowing the value of $k$ in advance. And $k$ could take any non-negative integer value. This computation cannot be achieved using bounded memory.
In [2], this is achieved non-deterministically by guessing the value of $k$, checking whether the heights of the two stacks are the same at the end of the process, and failing if they are not.
But we can achieve this deterministically by repeatedly pushing $\mathbf{q}^{\prime\mathsf{R}}$ until there are fewer than $s+t$ cells between the top of the pushdown stack and the top of the check stack (which can be tested by making the read write head move up while pushing padding symbols onto the pushdown stack), and at that stage pushing $\mathbf{p}^{\prime\mathsf{R}}$. At the end of this deterministic process we have $\mathbf{y}^{\delta}$ on the pushdown stack.
In the process described in [2], the cspda decides after processing each input symbol either to read another input symbol or to proceed to Stage 3, and look for no further input. If the machine moves to Stage 3 prematurely, then the end of the input word is never reached, and so that word is not accepted. If the machine looks for input after reading the end of the word, then it must also fail. The machine can only accept the word if it makes that transition at the right moment and finds that the contents of the two stacks are distinct. Clearly this part of the process of [2] is non-deterministic.
In our cspda, in order to maintain determinism, we compare the contents of the two stacks during our Stage 2 repeatedly, after we have processed each input symbol. We do this by popping symbols off the pushdown stack until we reach a point at which the two stacks differ; the first symbol of the pushdown stack that is not matched is then pushed back onto the stack. If the two stacks are distinct, then we move into the single accepting reading state $q_{2}$, but if they match we return to the single non-accepting reading (and entry) state $q_{1}$. In the event of no further input that state will be the final state. If there is further input then we read it, and then we copy across from the check stack to the pushdown stack to fill those cells at the top of the pushdown stack that we have just popped, so that both stacks have the same height as before, and contain the same contents as they did just before this checking process. We then proceed to process the new input symbol as already described above.
It is proved in [2] that the machine described there is a cspda that accepts an input word precisely when it is in $\mathsf{coWP}(G,A)$. It should be clear that our modified cspda accepts the same language as does the cspda of [2], and that our modifications ensure that properties (0), (1) and (2) hold for our cspda.
It remains to consider property (3). Since our machine never fails, we only need to check (3) (a), that for any words $u\in A^{*},v\in\mathsf{WP}(G,A)$ the configurations $C^{u}$ and $C^{uv}$ are equal. These configurations are determined by the state of the machine and the contents of the pushdown stack. At any stage the pushdown stack must contain the image of the string on the check stack under the action of the element represented by the input string that has been read so far. Since $u,uv$ represent the same group element, the pushdown stack contents in the two configurations must match, as required. And the state within each configuration must be either $q_{1}$ or $q_{2}$, depending on whether or not that single element is equal to the identity. $\Box$
5 Higman–Thompson groups
The Higman–Thompson groups $G_{n,r}$ groups were described by Higman in [8] as generalizations of one of Thompson’s finitely presented infinite simple groups, namely the group $V$ described in [15], which is isomorphic to $G_{2,1}$, and which first appeared in a handwritten manuscript of Richard J. Thompson circa 1965. They were proved to be co-indexed in [9] and the stronger result that they are co-context-free is proved in [12]. In this section we extend the proof of [9] to derive the following result.
Theorem 5.1.
The Higman–Thompson groups $G_{n,r}$ are special co-et0l-groups.
Note that, as finitely generated subgroups of the group $V\cong G_{2,1}$, the Thompson groups $F$ and $T$ are proved special co-et0l by the combination of Theorem 5.1 and Theorem 6.1 below.
There are several equivalent definitions of the groups $G_{n,r}$, and the one we use here is taken from [10, Section 1.10.11], and is also used in [9, Section 3]. Our proof that these groups are special co-et0l is based on the proof in [9, Theorem 1] that they are stack groups, and is similar to our proof of Theorem 4.1.
Fix integers $r\geq 1$ and $n\geq 2$, and let $Q:=\{q_{1},q_{2},\ldots,q_{r}\}$ and $\Sigma:=\{\sigma_{1},\sigma_{2},\ldots,\sigma_{n}\}$ be disjoint sets of cardinality $r$ and $n$ respectively. Define $\Omega$ to be the set of all infinite strings of the form $\omega=q_{i}\sigma_{i_{1}}\sigma_{i_{2}}\sigma_{i_{3}}\cdots$, where $q_{i}\in Q$ and $\sigma_{i_{j}}\in\Sigma$ for $j\geq 1$. A non-empty prefix of $\omega$ is a finite string (or word) $q_{i}\sigma_{i_{1}}\cdots\sigma_{i_{k}}$ for some $k\geq 0$. Roughly speaking $G_{n,r}$ is the group which acts on $\Omega$ by replacing non-empty prefixes; we give more detail below.
Observe that $Q\Sigma^{*}$ is the set of non-empty prefixes of strings in $\Omega$. We define a finite complete antichain in $\Omega$ to be a finite subset $B$ of $Q\Sigma^{*}$ so that each $\omega\in\Omega$ has precisely one element of $B$ as prefix. For example, when $r=1$ and $n=2$, $\{q_{1}\sigma_{1},q_{1}\sigma_{2}\sigma_{1},q_{1}\sigma_{2}\sigma_{2}\}$ is a finite complete antichain, but $\{q_{1}\sigma_{1},\,q_{1}\sigma_{1}\sigma_{1},\,\allowbreak q_{1}\sigma_{2}\sigma_{1},\,\allowbreak q_{1}\sigma_{2}\sigma_{2}\}$ is not.
A bijection $\varphi:B\to C$ between two finite complete antichains induces a permutation of $\Omega$, also called $\varphi$, defined by $\omega^{\varphi}:=b^{\varphi}w^{\prime}$, where $\omega=b\omega^{\prime}$ and $b$ is the unique prefix of $\omega$ that lies in $B$. The group $G_{n,r}$ is defined to be the group of all such permutations of $\Omega$. We also define $\mathbf{x}^{\varphi}:=b^{\varphi}\omega^{\prime}$ for any prefix $\mathbf{x}=b\omega^{\prime}$ of an element of $\Omega$ that has $b\in B$ as a prefix.
Proof of Theorem 5.1. This proof is very similar to that of Theorem 4.1 and the computation proceeds by making prefix replacements to a word stored on the pushdown stack. The main difference is that prefixes may be replaced by words of a different length. As in the earlier proof, there are just two reading states $q_{1}$ and $q_{2}$, which are respectively non-accepting and accepting, where the pushdown stack is empty in state $q_{1}$ and non-empty in state $q_{2}$. All non-reading states are non-accepting.
It is shown in [8] that $G_{n,r}$ is finitely presented. So we fix a finite inverse-closed set generating $A$ of $G_{n,r}$. Then there is a bound $K$ such that all finite complete antichains $B$ and $C$ occurring in the elements $(B,C,\varphi)$ of $A$ have length at most $K$.
In its first stage, the cspda with language $\mathsf{coWP}(G,A)$ puts an arbitrary word $\mathbf{x}$ in $Q\Sigma^{*}$ on the check stack, with its first symbol $q_{i}$ at the top, and then puts an arbitrary number of padding symbols on top of that. The read write head then moves to the bottom of the stacks, and the cspda enters its deterministic second stage in the non-accepting reading state $q_{1}$ with the pushdown stack empty. After reading the first letter (if any) of the input word $\alpha$, the word $\mathbf{x}$ is copied onto the pushdown stack. Thereafter, the cspda works by replacing $\mathbf{x}$ by $\mathbf{x}^{\beta}$, where $\beta$ is the prefix of $\alpha$ that has been read so far.
The cspda attempts to establish that the input word $\alpha$ does not fix some string in $\Omega$ with prefix $\mathbf{x}$, which would prove that it lies in $\mathsf{coWP}(G,A)$. To do that, it attempts to compute the image $\mathbf{x}^{\alpha}$ of $\mathbf{x}$ under $\alpha$, and then to check whether $\mathbf{x}^{\alpha}$ is unequal to $\mathbf{x}$, which would prove that $\alpha\in\mathsf{coWP}(G,A)$. If $\alpha$ does lie in $\mathsf{coWP}(G,A)$, then there exists a string in $\Omega$ that it does not fix and, provided that a sufficiently long prefix $\mathbf{x}$ of this string together with enough padding symbols are put on the check stack in the first stage, the cspda will succeed in showing that $w\not\in\mathsf{coWP}(G,A)$.
After reading the first input letter and copying $\mathbf{x}$ onto the pushdown stack, and also after reading all subsequent input letters $x=(B,C,\varphi)\in A$ and reconstituting the pushdown stack if necessary as explained below, the cspda looks for a prefix of the word $\mathbf{x}^{\beta}$ that is currently on the pushdown stack that lies in the set $B$. It can do this by examining the prefixes of $\mathbf{x}^{\beta}$ of length at most $K$. If $|\mathbf{x}^{\beta}|<K$ then it could fail to find such a prefix, in which case the computation fails. This means that the word $\mathbf{x}$ was not long enough. Otherwise a unique prefix $b\in B$ is identified, and this is replaced by $b^{\varphi}$ on the top of the pushdown stack. If $b^{\varphi}$ is longer than $b$, then this replacement might not be possible without going above the top of the check stack, in which case the computation fails. This occurs when not enough padding symbols were put at the top of the check stack. Otherwise, after making the replacement, the read write head is located at the top of the pushdown stack, and sees the first letter of the new word $\mathbf{x}^{\beta x}$. Since the processed prefix $\beta$ of $\alpha$ has now been replaced by $\beta x$, the word on the pushdown stack is now equal to $\mathbf{x}^{\beta x}$, as required.
As in the proof of Theorem 4.1, before entering a reading state, the cspda checks whether the words $\mathbf{x}^{\beta}$ and $\mathbf{x}$ on the stacks are equal. If they have different lengths or, equivalently, if the letter that the read write head sees on the check stack does not lie in the set $Q$, then the words cannot be equal, and the cspda moves immediately into the accepting state $q_{2}$, since there is an element of $\Omega$ that is not fixed. Otherwise, as in the proof of Theorem 4.1, the cspda checks the words for equality by emptying the pushdown stack of letters that are the same as the corresponding letters on the check stack. It then goes into state $q_{1}$ if the pushdown stack is empty, and into $q_{2}$ if it is non-empty. After reading the next input letter if any, if the pushdown stack is empty, or if the symbol currently at its top does not lie in $Q$, then the cspda reconstitutes the word $\mathbf{x}^{\beta}$ on the pushdown stack by copying letters from the check stack.
Checking that the required properties of a cspda hold for the automaton that we have defined is straightforward, and similar to the proof in Theorem 4.1. $\Box$
6 Closure properties of special co-et0l-groups
Our proofs of the closure properties in this section closely follow the proofs of corresponding properties for stack groups in [9, Sections 7 and 8].
Theorem 6.1.
The property of being a special co-et0l-group is independent of the choice of finite generating set. Furthermore, the class of special co-et0l-groups is closed under passing to finitely generated subgroups, finite direct products, and finite extensions.
Proof.
Let $G=\langle A\rangle$ be a special co-et0l-group and $H=\langle B\rangle\leq G$ with $B$ finite and closed under inversion. For each $b\in B$, we let $w_{b}$ be a specific word in $A^{*}$ with $b=_{G}w_{b}$. The special cspda $\mathcal{M}_{H}$ accepting $\mathsf{coWP}(H,B)$ is based on the special cspda $\mathcal{M}_{G}$ for $\mathsf{coWP}(G,A)$ but with input alphabet $B$ and with some extra states and non-reading moves. On reading a letter in $B$ in its deterministic second phase, the special cspda $\mathcal{M}_{H}$ behaves in the same way as $\mathcal{M}_{G}$ behaves on reading the word $w_{B}$. Note that by choosing $H=G$ this also proves the final assertion in the theorem.
Now let $G=H\times K$ with $H$ and $K$ special co-et0l-groups, where $\mathsf{coWP}(H,A_{H})$ and $\mathsf{coWP}(K,A_{K})$ are accepted by special cspda $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$. We construct a special cspda $\mathcal{M}_{G}$ accepting $\mathsf{coWP}(G,A_{H}\cup A_{K})$ as follows. It starts with an $\epsilon$-move into the initial state of either $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$, and thereafter behaves like $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ respectively. In the first case it ignores all input letters from $A_{K}$, and in the second case it ignores those from $A_{H}$.
[…]
The special cspda $\mathcal{M}_{G}$ accepting $\mathsf{coWP}(G,A)$ is based on $\mathcal{M}_{H}$, but its states are pairs $(q,t)$, where $q$ is a state of $\mathcal{M}_{H}$ and $t\in T$. The initial state is $(q_{0},1_{G})$, where $q_{0}$ is the initial state of $\mathcal{M}_{H}$. The first phase of $\mathcal{M}_{G}$ is the same as the first phase of $\mathcal{M}_{H}$, with the first component of the states changing as in $\mathcal{M}_{H}$ and the second component remaining $1_{G}$. In the second phase, when $\mathcal{M}_{G}$ reads the letter $a\in A$ in state $(q,t)$, it behaves as though $\mathcal{M}_{H}$ has read the word $w_{t,a}$ with the first state component changing accordingly, and the second state component changes to $t_{a}$. The cspda $\mathcal{M}_{G}$ accepts an input word $w$ if and only if, in the next reading state $(q,t)$ after reading $w$, either $t\neq 1_{G}$ or $q$ is an accept state of $\mathcal{M}_{H}$. ∎
Theorem 6.2.
The (standard) restricted wreath product $H\wr K$ of a special co-et0l group $H$ by a finitely generated virtually free top group $K$ is a co-et0l group.
Proof.
This is similar to the proof of the last part of [9, Proposition 9]. Let $H=\langle A_{H}\rangle$ be a special co-et0l-group and $\mathcal{M}_{H}$ a special cspda with $\mathcal{L}(\mathcal{M}_{H})=\mathsf{coWP}(H,A_{H})$. Let $K$ be finitely generated and virtually free. As in the proof of Theorem 3.1 above, we can choose a generating set $A_{K}=B\cup T^{\pm 1}$ of $K$ in which $B$ is the inverse closure of a free generating set of a free subgroup $F$ of finite index in $K$ and $T$ is a right transversal of $F$ in $K$ with the element in $F$ removed, and then every element of $K$ can be written uniquely as a word $wt$ with $w$ a freely reduced word in $B^{*}$ and $t\in T\cup\{\epsilon\}$. As we saw in the proof of Theorem 3.1, the word problem $\mathsf{WP}(K,A_{K})$ is the language of a pushdown automaton $\mathcal{P}_{K}$ that works by storing the normal form word $wt$ of the prefix of the input word read so far, and accepts by empty stack.
Let $G$ be the restricted wreath product $H\wr K$. Then $G=\langle A\rangle$ with $A=A_{H}\cup A_{K}$. (As usual, we assume that $A_{H}$ and $A_{K}$ are disjoint.) We shall describe a special cspda $\mathcal{M}_{G}$ with $\mathcal{L}(\mathcal{M}_{G})=\mathsf{coWP}(G,A)$. As in [9, Proposition 9] (and earlier results on closure under restricted wreath product), it works by non-deterministically choosing a component of the base group $H^{K}$ of $G$, and using $\mathcal{M}_{H}$ to attempt to prove the non-triviality of that component of an element of the base group. A word will be accepted by $\mathcal{M}_{G}$ if and only either the input word does not project onto the identity element of $K$, or $\mathcal{M}_{H}$ accepts the word representing the chosen component of the base group.
The initialisation of $\mathcal{M}_{G}$ consists of an initialisation of $\mathcal{M}_{H}$, followed by the writing of an arbitrary normal form word $wt$ on the check stack (with $t$ on top), on top of the initialised stack for $\mathcal{M}_{H}$. An arbitrary number of padding symbols p are then added to the check stack to help prevent the second extra condition of a special cspda being violated. This word $wt$ defines the component of the base group that we shall check for non-triviality using the initialisation of $\mathcal{M}_{H}$. The current state of $\mathcal{M}_{H}$ is stored at all times as a component of the state of $\mathcal{M}_{G}$.
After reading the first input symbol $x$, the read write head of $\mathcal{M}_{G}$ ascends to the top of the check stack of $\mathcal{M}_{H}$ (while pushing padding symbols to the pushdown stack) and then copies the word $wt$ from the check stack to the pushdown stack. Then it passes $x$ to $\mathcal{P}_{K}$ for processing using its pushdown stack (which is the upper part of the pushdown stack of $\mathcal{M}_{G}$).
Thereafter, the automaton $\mathcal{P}_{K}$ processes input symbols $x$ in this way. But whenever the pushdown stack of $\mathcal{P}_{K}$ is empty and an input symbol $x\in B$ is read, the read write head descends while removing padding symbols from the pushdown stack, and $x$ is processed by $\mathcal{M}_{H}$ using its current stored state. After $\mathcal{M}_{H}$ has processed $x$, the read write head moves back up and, unless $\mathcal{M}_{G}$ is in its entry configuration, control is passed back to $\mathcal{P}_{K}$ for processing the next input letter.
There is one further feature that is necessary for $\mathcal{M}_{G}$ to ignore the word problem of $G$. Suppose that at some reading configuration $\mathcal{M}_{H}$ is in its entry state, and the height of the pushdown stack of $\mathcal{P}_{K}$ is the same as the height of the check stack minus the padding symbols at the top of the check stack. Then $\mathcal{M}_{G}$ checks whether it should be in its entry configuration by checking whether the symbols read on the two stacks are the same and, if so, popping from the pushdown stack. It repeats this process while the symbols read are the same. If at some stage before it reaches the bottom of the pushdown stack of $\mathcal{P}_{K}$ the read symbols are not the same, then the input word does not project onto the identity element of $K$, so $\mathcal{M}_{G}$ is not in its entry configuration, and any popped symbols are replaced. If, however, this process results in the pushdown stack of $\mathcal{P}_{K}$ being emptied, then the read write head descends to the bottom of the stack of $\mathcal{M}_{G}$, and $\mathcal{M}_{G}$ returns to its entry configuration. After reading the next letter (if any), the check stack is reconstituted before processing that letter.
The accept states of $\mathcal{M}_{G}$ are all reading states other than the entry states. It is now straightforward to verify that the cspda that we have described is special. Property (3) follows from the assumption that $\mathcal{M}_{H}$ has this property. ∎
Theorem 6.3.
The class of special co-et0l-groups is closed under taking free products with finitely many free factors.
Proof.
It is enough to prove that the free product $G=H*K$ of two special co-et0l-groups $H$ and $K$ is a special co-et0l-group. Let $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$ be special cspda accepting $\mathsf{coWP}(H,A_{H})$ and $\mathsf{coWP}(K,A_{K})$ respectively. We shall define a special cspda $\mathcal{M}_{G}$ accepting $\mathsf{coWP}(G,A)$ with $A=A_{H}\cup A_{K}$ (which we assume to be a disjoint union). The cspda $\mathcal{M}_{G}$ includes disjoint copies of $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$. Beyond the single initial state of $\mathcal{M}_{G}$, each other state of $\mathcal{M}_{G}$ is naturally associated with a state of either $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$. However, since $\mathcal{M}_{G}$ needs to do some additional processing while remembering its current state of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$, that current state will be stored only as the first component of a state of $\mathcal{M}_{G}$, which will also contain some further information. We denote by $\Delta_{H}$ and $E_{H}$ the stack alphabet and set of entry states for $\mathcal{M}_{H}$, and by $\Delta_{K}$ and $E_{K}$ the stack alphabet and set of entry states for $\mathcal{M}_{K}$.
In the first phase (i.e. the initialisation) of $\mathcal{M}_{G}$ an arbitrary string $\alpha_{1}\beta_{1},\cdots\alpha_{k}\beta_{k}$ for some $k\geq 1$ is put onto the check stack (with $\alpha_{1}$ at the bottom), with substrings $\alpha_{i}$ and $\beta_{i}$ over $\Delta_{H}\cup E_{H}$ and $\Delta_{K}\cup E_{K}$ respectively, which are all non-empty for $i=1,\ldots,k$, except possibly for $\alpha_{1}$ and/or $\beta_{k}$ (since the string could start or end with a substring over either $\Delta_{H}\cup E_{H}$ or $\Delta_{K}\cup E_{K}$). Each non-empty string $\alpha_{i}$ or $\beta_{i}$ consists of a string over $\Delta_{H}$ or $\Delta_{K}$ that is valid for the initialisation of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ followed by the entry state of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ that results from that initialisation. The check stack of $\mathcal{M}_{G}$ that we create during this initialisation is seen to be a sequence of check substacks that are initialised stacks of $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$. We let $q_{\mathsf{ent}}$ be the entry state of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ that is at the top of the lowest of those check substacks (corresponding to the first non-empty substack); we store $q_{\mathsf{ent}}$ as a component of the state of $\mathcal{M}_{G}$.
After the check stack has been set, the read write head of $\mathcal{M}_{G}$ descends to the bottom of the pair of stacks. Then the first component of the entry state of $\mathcal{M}_{G}$ is set to be the entry state $q_{\mathsf{ent}}$ that was recorded at the top of the bottom check substack. This concludes the first (initialisation) phase.
At all times during the second phase, exactly one of the subgroups $H$ and $K$ will be the current subgroup. When the current subgroup is $H$, the cspda $\mathcal{M}_{G}$ will have a state of $\mathcal{M}_{H}$ as its first component, and its read write head will be scanning a check substack of $\mathcal{M}_{H}$, which we call the current check stack, and similarly for $K$. The general idea is that, when the current subgroup is $H$ and input symbols from $A_{H}$ are being read, the cspda behaves in the same way as $\mathcal{M}_{H}$, using the current check substack. We call such a sequence of moves of $\mathcal{M}_{G}$ a sub-computation in $\mathcal{M}_{H}$ or in $\mathcal{M}_{K}$. As we shall see later, the cspda needs to remember the entry state of the current sub-computation, and we can store this as the second component of the current state. At the start of the second phase, the read write head is at the bottom of the stacks of $G$, the current subgroup is $H$ or $K$, depending on whether the first check substack contains $\alpha_{1}$ or $\beta_{1}$, and the current state of $\mathcal{M}_{G}$ (i.e. the entry state of this computation) is $(q_{\mathsf{ent}},q_{\mathsf{ent}})$.
We need to describe how $\mathcal{M}_{G}$ behaves when it reads a letter $x$ from the input tape. Suppose first that the current subgroup is $H$. Now suppose also that $x\in A_{K}$, so $x$ is not in the current subgroup. In this case, $\mathcal{M}_{G}$ moves to the top of that current check substack while pushing padding symbols p onto the pushdown stack. At the top of this check substack, it records the current state of $\mathcal{M}_{H}$ on the pushdown stack (in the same position as the current entry state is stored on the check substack). If there are no further check substacks above the current one, then the initialisation stage of $\mathcal{M}_{G}$ was too short and $\mathcal{M}_{G}$ fails; we say that the initialisation was inadequate. Otherwise, the read write head moves up into the next check substack, which will be for the group $K$. The read write head first moves to the top of this check substack (while pushing padding symbols onto the pushdown stack) to retrieve the entry state $q_{\mathsf{ent}}$ of this initialisation of $\mathcal{M}_{K}$. Then it descends again to the bottom of this check substack (popping from the pushdown stack as it does so) and behaves as though $\mathcal{M}_{K}$ had read the input symbol $x$ while in entry state $q_{\mathsf{ent}}$. It changes to the state $(q,q_{\mathsf{ent}})$, where $q$ is the state into which a sub-computation on $\mathcal{M}_{K}$ would transition from $q_{\mathsf{ent}}$, given input $x$ and the current stack symbols, and pushes to the pushdown stack, accordingly.
Otherwise we have $x\in A_{H}$. As explained earlier, $\mathcal{M}_{G}$ now behaves in the same way as $\mathcal{M}_{H}$, using the currently scanned check substack as its check stack. Where $q_{\mathsf{ent}}$ is the entry state of this sub-computation, and $q$ is the state that $\mathcal{M}_{H}$ should enter after reading $x$, $\mathcal{M}_{G}$ will now enter the state $(q,q_{\mathsf{ent}})$.
So far we have assumed that the current subgroup is $H$. If the current subgroup is $K$ then the machine operates in just the same way, but with the roles of $H$ and $K$ in the description above reversed.
If, after completing a reading move and any subsequent non-reading moves of a sub-computation in the subgroup $H$, the cspda is in the entry configuration $q_{\mathsf{ent}}$ of the current sub-computation within $\mathcal{M}_{H}$ (that is, if the current state is $(q_{\mathsf{ent}},q_{\mathsf{ent}})$), but not in the entry configuration of $\mathcal{M}_{G}$ (so the read write head is at the bottom of its current check substack, but not at the bottom of the check stack of $\mathcal{M}_{G}$), then the current subgroup changes from $H$ to $K$, and the read write head descends into the check substack immediately below its current position, which will be for the group $K$. This lower check substack has recorded the configuration of a sub-computation within $\mathcal{M}_{K}$ that had been interrupted when a letter from $A_{H}$ was read. The read write head retrieves the entry state and the state of that sub-computation at the point of interruption from the tops of the check substack and pushdown stack, respectively, and then it descends and removes any padding symbols p from the top of the pushdown stack. After doing that, the configurations of both $\mathcal{M}_{G}$ and of the sub-computation in $\mathcal{M}_{K}$ are the same as they were at the point of interruption. Note that, if the next input letter is from $A_{H}$ then $\mathcal{M}_{G}$ immediately goes up again and resumes the current sub-computation in $\mathcal{M}_{H}$, but this temporary change of current subgroup and descent into the lower check substack is necessary to ensure that $\mathcal{M}_{G}$ ignores the word problem of $G$.
Reversing the roles of $H$ and $K$ in the paragraph above deals with the case where the cspda enters an entry configuration of the current sub-computation within $\mathcal{M}_{K}$ but is not in an entry configuration of $\mathcal{M}_{G}$. That completes the description of the cspda $\mathcal{M}_{G}$.
We shall show next that, so long as $\mathcal{M}_{H}$ and $\mathcal{M}_{K}$ are special then so is $\mathcal{M}_{G}$. It should be clear that $\mathcal{M}_{G}$ satisfies the additional properties (0), (1) and (2). It remains to verify property (3), that $\mathcal{M}_{G}$ ignores the word problem of $G$. So let $C$ be an entry configuration of $\mathcal{M}_{G}$ and let $uv\in A^{*}$ with $v=_{G}1$, and assume also that $C^{uv}$ does not fail. We need to show that $C^{u}=C^{uv}$.
Suppose first that the letters of $v$ are all in the same alphabet $A_{H}$ or $A_{K}$ we may assume without loss that $v\in A_{H}^{*}$. Then $v=_{H}1$, and the claim follows essentially from the assumption that $\mathcal{M}_{H}$ ignores the word problem of $H$, because $\mathcal{M}_{G}$ is carrying out a sub-computation in $\mathcal{M}_{H}$ while reading $v$. There are some complications however.
If the current subgroup of the configuration $C^{u}$ is $K$ then, on reading the first letter of $v$, the current subgroup changes to $H$. Then, after reading $v$, $\mathcal{M}_{H}$ will be in an entry configuration of $\mathcal{M}_{H}$, and the read write head will descend into the check substack of $\mathcal{M}_{K}$, retrieving the reading state $M_{K}$ was left at and possibly removing some padding symbols, resulting in $C^{uv}=C^{u}$.
It is also possible that, after reading a letter $x$ of $v$ that is not the final letter of $v$, the current sub-computation within $\mathcal{M}_{H}$ is in an entry configuration of $\mathcal{M}_{H}$, in which case (as we have explained above) the stack read write head of $\mathcal{M}_{G}$ descends into the check substack below. Since $x$ is not the last letter of $v$, then on reading the following letter of $v$, the read write head moves back up and the sub-computation resumes in $\mathcal{M}_{H}$.
If $v$ contains letters from both $A_{H}$ and $A_{K}$, then there must be some non-trivial subword $v^{\prime}$ of $v$ that contains letters only from $A_{H}$ or $A_{K}$ with $v^{\prime}=_{H}1$ or $v^{\prime}=_{K}1$ and then the claim follows from the previous paragraphs and induction on the number of components of $v$ (where a component is a maximal non-empty subword of $v$ within either $A_{H}^{*}$ or $A_{K}^{*}$).
It remains to prove that $\mathcal{L}(\mathcal{M}_{G})=\mathsf{coWP}(G,A)$. Since the entry states of $\mathcal{G}$ are not accept states, and $\mathcal{M}_{G}$ ignores the word problem of $G$, no word $w\in A^{*}$ with $w=_{G}1$ can be accepted by $\mathcal{M}_{G}$, and so we have $\mathcal{L}(\mathcal{M}_{G})\subseteq\mathsf{coWP}(G,A)$.
Conversely, suppose that $w\in\mathsf{coWP}(G,A)$, and write $w=w_{1}\cdots w_{k}$ as a sequence of components. We shall prove by induction on the number $k$ of components that $w\in\mathcal{L}(\mathcal{M}_{G})$. In fact we shall use Lemma 2.1 to prove the stronger statement that, for any $n\in{\mathbb{Z}}$ with $n>0$, there is an entry configuration $C$ of $\mathcal{M}_{G}$ such that
  • (i)
    $C^{w}$ is an accept state of $\mathcal{M}_{G}$; and
  • (ii)
    for each entry configuration $C_{i}$ of $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$ that is defined by one of the check substacks in the initialisation of $\mathcal{M}_{G}$ that results in $C$, and any word $v$ over $A_{H}$ or $A_{K}$ of length at most $n$, $C_{i}^{v}$ does not fail.
If $k=1$ then $w$ is a word over $A_{H}$ or $A_{K}$ in either $\mathsf{coWP}(H,A)$ or $\mathsf{coWP}(K,A)$, and so $\mathcal{M}_{G}$ operates as either $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$, whose languages are both within $\mathcal{L}(\mathcal{M}_{G})$. The stronger statement in the preceding paragraph follows from Lemma 2.1 applied to $\mathcal{M}_{H}$ or $\mathcal{M}_{K}$. So suppose that $k>1$.
If each component of $w$ lies in either $\mathsf{coWP}(H,A_{H})$ or $\mathsf{coWP}(K,A_{K})$, then each lies in either $\mathcal{L}(\mathcal{M}_{H})$ or $\mathcal{L}(\mathcal{M}_{K})$, and there are entry configurations of $\mathcal{M}_{H}$ and of $\mathcal{M}_{K}$ that result in these components being accepted and also, by Lemma 2.1, do not fail on input of any word over $A_{H}$ or $A_{K}$ of length at most $n$. By initialising $\mathcal{M}_{G}$ with the corresponding sequence of initialisations of $\mathcal{M}_{H}$ and of $\mathcal{M}_{K}$, we ensure that each component of $w$ is processed by a sub-computation in a separate substack and, after reading $w$, the cspda $\mathcal{M}_{G}$ is in the accept state arising from the sub-computation of $w_{k}$. So $w\in\mathcal{L}(\mathcal{M}_{G})$ with the conditions of the stronger assertion satisfied.
Otherwise choose $i$ such that the component $w_{i}$ of $w$ lies within the word problem of one of $H$ or $K$, and define $u:=w_{1}\cdots w_{i-1}$, $u^{\prime}:=w_{i+1}\cdots w_{k}$, $v:=w_{i}$; if $i=1$ or $i=k$ then $u$ or $u^{\prime}$ is defined to be empty. Since, by assumption, we have $w=uvu^{\prime}\in\mathsf{coWP}(G,A)$ and $v=_{G}1$, we have $uu^{\prime}\in\mathsf{coWP}(G,A)$ and $v\in\mathsf{WP}(G,A)$. By inductive hypothesis, $uu^{\prime}$, which has at most $k-1$ components, is within $\mathcal{L}(\mathcal{M}_{G})$, so there is an entry configuration $C$ of $\mathcal{M}_{G}$ for which $C^{uu^{\prime}}$ is an accept configuration. Furthermore, for any $n>0$, we can choose $C$ to satisfy Condition (ii) above, and in any case we choose it to satisfy this condition with $n$ equal to the length of $w$.
We claim that we can choose $C$ such that $C^{uv}$ does not fail and such that $C$ satisfies Condition (ii) above. By Condition (ii) applied to our initial choice of $C$ such that $C^{uu^{\prime}}$ is an accept state, the only way that $C^{uv}$ could fail would be if $C$ was an inadequate initialisation for reading $C^{uv}$, in which case we would require a single extra check substack to read the subword $v$ of $uv$. So we simply replace $C$ by the same configuration but with the required additional substack on top that satisfies Condition (ii) for the chosen $n$ assumed to be at least equal to the length of $w$. (We can assume that $H$ and $K$ are both non-trivial, and so $\mathcal{L}(\mathcal{M}_{H})$ and $\mathcal{L}(\mathcal{M}_{K})$ are non-empty, and we can apply Lemma 2.1 to choose this new substack.)
Now, since $\mathcal{M}_{G}$ ignores the word problem (i.e. we have Property (3)) and $C^{uv}$ does not fail, we have $C^{w}=C^{uvu^{\prime}}=C^{uu^{\prime}}$ is an accept state, and $C$ satisfies Condition (ii) above as required. ∎